OpenStudy (anonymous):

find the second derivative of (4-2x^2)^10

OpenStudy (anonymous):

The answer is apparently 80(4-2x^2)(19x^2-2) and I am NOT getting it...

OpenStudy (anonymous):

y' = 10(4-2x^2)^9 * (-4x)

OpenStudy (anonymous):

did you get that

OpenStudy (anonymous):

Yep! I got that exactly!

OpenStudy (anonymous):

\[\frac{d}{dx}(4-2x^{2})^{10} = 10*(4-2x^{2})^{9}*(-4x) = -40x*(4-2x^{2})^{9}\] // chain rule \[\frac{d}{dx}(-40x*(4-2x^{2})^{9}) = -40*(4-2x^{2})^{9} + (-40x*9*(4-2x^{2})^{8}*(-4x))\] \[= [-40*(4-2x^{2})^{8}]*[(4-2x^{2})+9*(x*-4x)]\] \[=[ -40*(4-2x^{2})^{8}]*[(4-2x^{2}-36x^{2}]\] \[= [ -40*(4-2x^{2})^{8}]*[(4-38x^{2})]\]

OpenStudy (anonymous):

next use the product rule so y'' = 10(4-2x^2)*(-4) + (10(-4x))(-4x)

OpenStudy (anonymous):

sorry hold on i forgot the x^9

OpenStudy (anonymous):

y' = 10(4-2x^2)^9 * (-4x) y'' = (10(4-2x^2)^9) * (-4) + (10*8(4-2x^2)^8)*-4x*-4x

OpenStudy (anonymous):

then simplify

OpenStudy (anonymous):

AHA! Thanks soooo much! I'm reviewing right now, so I'm trying to troubleshoot.

OpenStudy (anonymous):

your welcome, guess @tomo already had it covered for ya though

OpenStudy (anonymous):

Okay wait a sec... where'd the two -4x's come from?

OpenStudy (anonymous):

from takin the second half of the product rule y' = 10(4-2x^2)^9 * (-4x) when you do d/dx(10(4-2x^2)^9) * -4x in the d/dx you have another chain rule

OpenStudy (anonymous):

wouldnt it be -4x(y'(10(4-2x^2)^9)+10(4-2x^2)^9(y'(-4x)

OpenStudy (anonymous):

well the first half of the product rule the way you are doing it.

OpenStudy (anonymous):

-4x(y'(10(4-2x^2)^9) this part d/dx(10(4-2x^2)^9) you have to use a chain rule and so another -4x pops out

OpenStudy (anonymous):

Ack! I thought I missed one of the little buggers! Okay, I've got it now! Thanks again for the help!

OpenStudy (anonymous):

yep just simplify it and make sure its right. its easy to mess up these things

OpenStudy (anonymous):

that may not be right and i dont think the answer you said it was was right either. the correct answer is 40 (4-2 x^2)^8 (38 x^2-4)

OpenStudy (anonymous):

thats what i got using wolfram alpha

OpenStudy (anonymous):

The answers I'm given are never on par with wolfram alpha... would yo mind would checking me here? This is from using -40x(4-2x^2)^9 (-40x(9(4-2x^2)^8(-4x))+(10(4-2x^2)^9*(-40)

OpenStudy (anonymous):

did you forget a 10 in the first part

OpenStudy (anonymous):

since there is one in the second

OpenStudy (anonymous):

so then: (-40x(90(4-2x^2)^8(-4x))+(10(4-2x^2)^9*(-40)

OpenStudy (anonymous):

caught a mistake in my answer y'' = (10(4-2x^2)^9) * (-4) + (10*8(4-2x^2)^8)*-4x*-4x shoulb be y'' = (10(4-2x^2)^9) * (-4) + (10*9(4-2x^2)^8)*-4x*-4x ^

OpenStudy (anonymous):

So my answer is correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ugh... I hate this stuff. What am I doing wrong?!

OpenStudy (anonymous):

one step at a time y' = 10(4-2x^2)^9 * (-4x) product rule (d/dx)1st * 2nd = d/dx 2nd *first for d/dx 1st*2nd d/dx(10(4-2x^2)^9) = 9*10(4-2x^2)^8*(-4x) * -4x second for d/dx 2nd * 1st d/dx(-4x) = -4 so te wholpart is -4(10(4-2x^2)^9) then add the two parts 9*10(4-2x^2)^8*(-4x) * -4x -4(10(4-2x^2)^9)

OpenStudy (anonymous):

AHA! Okay I've got it now! Thanks again for all the help!

OpenStudy (anonymous):

no prob man